A man travel from his home to office at 4km/hr and reaches his office 20 minute late if the speed had been 6km/hr he would have 10 minute early.find the distance from his home to office

Grade:12

1 Answers

Susmita
5 years ago
Say the time within which he should reach is t hr.
Remember the formula x=vt.
In the 1st case,he reaches 20mint=20/60 hr=1/3 hr late.
So x=4(t+1/3)
Or,3x=4(3t+1)
Or,3x=12t+4
In the second case he take 10 mintutes=10/60 hr=1/6 hr less to reach.
So x=6(t-1/6)
Or,6x=36t-6
Multiply the 1st equation by 3 and then subtract from equation 2.You will get
Please approve the answer if you are helped
-3x=-18
Or,x=6
x=6km